OpenStudy (anonymous):

Please Help the pep club is going to sell bouquets of flowers during the homecoming game. they represent their revenue R and costs C with the following equations. R = 5.50x C=250+1.25x x is the number of bouquets What is the Y intercept of the line for eatch equation? What information does it represent in this context?

OpenStudy (anonymous):

R - is y for this equation R = 5.50x C- is y for this equation C = 250+1.25x and to find the y-intercept x must be equal to 0 so For R = 5.50x , y-intercept is 0 For C = 250 + 1.25x , y-intercept = 250 Correct me if I'm wrong :)